# A comparison between best-fit eccentricity definitions and the standardized definition of eccentricity

### v1

@@ I Introduction
Consequently, the GW community has undertaken significant efforts to develop tools and frameworks for detecting and analyzing GW data from such binaries on eccentric orbits. Several groups have been actively developing gravitational waveform models for binaries in eccentric orbit using semi-analytic formalism like Effective-One-Body (EOB) and Post-Newtonian approximation as well as employing fully Numerical Relativity (NR) simulations Warburton et al. (2012); Osburn et al. (2016); Cao and Han (2017); Liu et al. (2020); Ramos-Buades et al. (2022a); Nagar et al. (2021); Islam et al. (2021); Liu et al. (2022); Memmesheimer et al. (2004); Huerta et al. (2014); Tanay et al. (2016); Cho et al. (2022); Moore et al. (2018); Moore and Yunes (2019); Van De Meent and Warburton (2018); Chua et al. (2021); Hughes et al. (2021); Katz et al. (2021); Lynch et al. (2022); Klein (2021). Several works also have demonstrated the need to include eccentricity in the waveform models by showing potential bias in estimated parameters and spurious violation of General Relativity due to neglecting eccentricity Favata et al. (2022); O’Shea and Kumar (2023); Saini et al. (2022, 2023); Favata (2014); Shaikh et al. (2024).
@@
Modelling waveforms from compact binaries in eccentric orbits has one challenging aspect. Unlike Newtonian orbits, the orbit of a binary in a strong gravity regime can not be described by a closed ellipse – the orbit goes through pericenter precession. In addition, due to GW radiation, the size of the orbit and the eccentricity reduce towards merger time Zevin et al. (2019). There is no unique definition of eccentricity in General Relativity and, therefore, to capture these features of a binary in an eccentric orbit, different waveform models adopt custom internal definitions as per the convenience of the framework of the model. This leads to an unwanted ambiguity regarding the eccentricity parameter of the waveform models Knee et al. (2022). Such ambiguity could be reflected in the posteriors of parameter inference from GW strain data that include eccentricity, and would impede our understanding of astrophysical populations among research groups choosing different eccentric waveform models.
@@
To avoid such undesired ambiguity, there has been an effort to standardize the definition of eccentricity based on a gauge-independent quantity like the GW modes at infinity Ramos-Buades et al. (2020); Islam et al. (2021); Bonino et al. (2023); Ramos-Buades et al. (2022b); Shaikh et al. (2023). Since such a definition uses only the information in the GW modes, it removes the model dependency, i.e. it is independent of the way the GW modes are generated. This definition makes use of the oscillatory nature of the frequency and amplitude evolution of the eccentric GWs–it determines the locations of the local maxima and minima obtained from waveform quantities like the frequency or the amplitude. Consequently, the computation requires the waveform modes to include a minimum (typically $\gtrsim 5$) number of orbits to reliably determine the eccentricity, thus far the definition has been examined primarily for waveforms in which the $(2,2)$ multipole is dominant. See Ref. Shaikh et al. (2023) for more details on the available strategies to implement a standardized definition of eccentricity, which we refer to as $e_{\text{gw}}$. Recently, Patterson et al. (2024) proposed an eccentricity identification, also restricted to the $(2,2)$-mode, from an harmonic decomposition of the waveform and Boschini et al. (2025) developed a gauge independent estimation method from catastrophe theory.
@@
Although advances in NR formalism and the availability of large computational resources have enabled NR groups to simulate increasingly long compact binary coalescences (CBCs), including eccentric ones, many NR simulations still lack a sufficient number of orbits to reliably apply the standardized definition of eccentricity. For such shorter NR simulations, other definitions of eccentricity may be used to assign an eccentricity value to these simulations. One commonly used definition of eccentricity involves the minimization of a Loss Function (LF) by adjusting the parameters of a chosen waveform model to best match the given NR waveform. The mismatch $\mathcal{M}$ Flanagan and Hughes (1998)111or equivalently its counterpart to maximize, the overlap or faithfulness e.g. in Hinder et al. (2018), is commonly used in such procedure since this scalar quantity defines the optimal matched filter in the GW detection problem. The mismatch can either play the role of the LF to minimize (e.g. Joshi et al., 2023) or can be used as a validation metric after minimization of another LF. For instance, the authors of Habib and Huerta (2019) computed their LFs from the evolution of the dimensionless quantity $M\omega$, with $M$ the BBH total mass and $\omega$ the mean orbital frequency, and then calculated the mismatch to quantify the accuracy of the characterization of the eccentric NR simulation at hand.
@@
Alternatively, one can choose the $L_{2}$-norm, the distance between vectors of strain data in time domain as the LF to compare a model (waveform, surrogate model to train) from a set of NR simulations Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024); Rink et al. (2024). We will thus also include this $L_{2}$-norm in our analysis.
@@ III.1 NR simulations and model waveform
For this study, we use $10$ eccentric simulations, as listed in Table 2, from the Simulating eXtreme Spacetimes (SXS) catalog (Boyle et al., 2019; SXS Collaboration, ; SXS, ). We sort them in this table by $e_{\text{gw}}$ at $18$Hz from the measurements of Fig. 1. We use the implementation in the Python library gw_eccentricity  [^2] https://github.com/vijayvarma392/gw_eccentricity from Shaikh et al. (2023) to compute $e_{\text{gw}}$ in this work. We explain the chosen method later in Section III.2.
@@
To assign $e^{\text{best-fit}}$ to these simulations, we use the publicly available eccentric waveform model TEOBResumS-Dalí  [^3] https://bitbucket.org/teobresums/teobresums/. We have used the git commit 0f19532a9a1012b556bd4c72da10a1317d0e8751. To compute $e_{\text{gw}}$ and the LF $\mathcal{E}$, we will make use of the dominant (2, 2) mode only and output physical units strain from TEOBResumS-Dalí, In addition, since TEOBResumS-Dalí allows to define the reference frequency as the waveform frequency apastron, periastron or the mean of the two, we choose the latter definition as in Knee et al. (2022). The case including higher modes has notably been investigated by e.g. Habib and Huerta (2019); Bonino et al. (2023), and is beyond the scope of our study. It should also be kept in mind that the definition of the eccentric initial conditions are sensitive to changes in input parameters $e_{0},f_{0}\sim\mathcal{O}(10^{-6})$ Nagar and Rettegno (2021).
@@ III.2 Measuring $e_{\text{gw}}$
Figure 1: Left:Frequency evolution of eccentricity $e_{\text{gw}}$ in the simulations from Table. 2. Right:: $e_{\text{gw}}$ ratio between the ResidualAmplitude and Amplitude method from gw_eccentricity Shaikh et al. (2023). The former is more robust to the small eccentric NR strains in our set, justifying our preference. We resort to physical units by choosing a total binary mass $M=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
@@ III.3.2 Loss Functions
    We will note, for simplicity, $\mathcal{E}_{\mathpzc{h}_{22}}(f_{0},e_{0})=\|\delta\mathpzc{h}_{22}\|$, also emphasizing that this scalar is a distance in time-domain functions and not a correlation measure of Fourier space vectors as the mismatch. This LF is at the core of works such as Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024); Rink et al. (2024), notably to fit surrogate parametric models to reference NR simulations.
@@ V Conclusion and remarks
In addition, the recent study Bonino et al. (2024) would corroborate this prescription on the mismatch, although not in the context of numerical searches as in our work. The authors, after assigning eccentricities to their own NR simulations set through a mapping procedure of the initial conditions, find $\mathcal{M}\sim 10^{-3}-10^{-2}$ (Figure 12 of current version) as a robustness metric rather than a LF. We argue that $C_{\lambda}$ is merely an alternate LF to deal with both parametric model limitations and mismatch requirements. A simple to implement LF $C_{\lambda}$ will be particularly useful in the context of fitting surrogates to reference strains, among the practical applications that motivated our study. We leave new developments on fitting procedures invoking $e_{\text{gw}}$ in addition to classic LFs, with numerical searches or other strategies as in Bonino et al. (2024), to future studies.
@@ A.2 LF minimization results
Table 2: Characterization of SXS simulations with TEOBResumS-Dalí using different LFs, along with standardized $e_{\text{gw}}$ measurements. Blue and red entries correspond to Figure 11. The leftmost column indicates the chosen LF, and the "Values" are computed from the TEOBResumS-Dalí waveform that minimizes this LF. From left to right, the SXS:BBH:XXXX IDs are sorted by increasing $e_{\text{gw}}$ from Table 2. Cases where the best-fit model has $\delta e_{\text{gw}}\sim 100\%$ correspond to a preferred quasi-circular model by that LF, which is the case when $LF\sim\mathcal{M}$ for the small eccentricity cases.

### v2

@@ I Introduction
Consequently, the GW community has undertaken significant efforts to develop tools and frameworks for detecting and analyzing GW data from such binaries on eccentric orbits. Several groups have been actively developing gravitational waveform models for binaries in eccentric orbit using semi-analytic formalism like Effective-One-Body (EOB) and Post-Newtonian approximation as well as employing fully Numerical Relativity (NR) simulations Warburton et al. (2012); Osburn et al. (2016); Cao and Han (2017); Liu et al. (2020); Ramos-Buades et al. (2022a); Nagar et al. (2021); Islam et al. (2021); Liu et al. (2022); Memmesheimer et al. (2004); Huerta et al. (2014); Tanay et al. (2016); Cho et al. (2022); Moore et al. (2018); Moore and Yunes (2019); Van De Meent and Warburton (2018); Chua et al. (2021); Hughes et al. (2021); Katz et al. (2021); Lynch et al. (2022); Klein (2021); Wang et al. (2025); Shi et al. (2025); Phukon et al. (2025); Sridhar et al. (2024); Gamboa et al. (2024a, b); Morras et al. (2025); Planas et al. (2025). , Several works demonstrated the need to include eccentricity by showing potential bias in estimated parameters and spurious violation of General Relativity due to neglecting eccentricity Favata et al. (2022); O’Shea and Kumar (2023); Saini et al. (2022, 2023); Favata (2014); Shaikh et al. (2024); Das et al. (2024); Kumar et al. (2025). The inclusion of other physical effects of eccentric binaries, such as the argument of periapsis (e.g. Clarke et al., 2022), is also being discussed.
@@
Modelling waveforms from compact binaries in eccentric orbits has one challenging aspect. Unlike Newtonian orbits, the binary orbit in a strong gravity regime can not be described by a closed ellipse – the orbit goes through pericenter precession. In addition, due to GW radiation, the size of the orbit and the eccentricity reduce towards merger time Zevin et al. (2019). There is no unique definition of eccentricity in General Relativity and, therefore, to capture features of a binary in an eccentric orbit, different waveform models adopt internal definitions as per the convenience of the model framework. This leads to an unwanted ambiguity regarding the eccentricity parameter of waveform models Knee et al. (2022). Such ambiguity could be reflected in the posteriors of parameter inference from GW strain data that include eccentricity, and would impede our understanding of astrophysical populations among research groups choosing different eccentric waveform models.
@@
To avoid such ambiguity, there has been efforts to standardize the definition of eccentricity based on a gauge-independent quantity like the GW modes at infinity Ramos-Buades et al. (2020); Islam et al. (2021); Bonino et al. (2023); Ramos-Buades et al. (2022b); Shaikh et al. (2023). Since such a definition uses only the information in the GW modes, it removes the model dependency, i.e. it is independent of the way the GW modes are generated. This definition makes use of the oscillatory nature of the frequency and amplitude evolution of the eccentric GWs–it determines the local maxima and minima from waveform quantities like the frequency or the amplitude. Consequently, the computation requires a minimum (typically $\gtrsim 5$) number of orbits in the waveform to reliably estimate eccentricity. Thus far, the definition has been examined primarily for waveforms in which the $(2,2)$ multipole is dominant. See Ref. Shaikh et al. (2023) for more details on strategies to implement a standardized definition of eccentricity, which we refer to as $e_{\text{gw}}$. Recently, Patterson et al. (2025) proposed an eccentricity identification, also restricted to the $(2,2)$-mode, from an harmonic decomposition of the waveform and Boschini et al. (2025) developed a gauge independent estimation method from catastrophe theory. The authors of Islam and Venumadhav (2025) propose a publicly available framework for smooth, monotonic eccentricity estimates invoking universal modulation features of non-precessing binaries Islam (2024a); Islam and Venumadhav (2024).
@@
Although advances in NR formalism and large computational resources have enabled NR groups to simulate increasingly long compact binary coalescences (CBCs), including eccentric ones, many NR simulations lack a sufficient number of orbits to reliably apply the standardized definition of eccentricity. For such shorter NR simulations, other definitions of eccentricity may be used to assign an eccentricity value to these simulations. One commonly used definition of eccentricity involves the minimization of a Loss Function (LF) by adjusting the parameters of a chosen waveform model to best match the given NR waveform. The mismatch $\mathcal{M}$ Flanagan and Hughes (1998)111or equivalently its counterpart to maximize, the overlap or faithfulness e.g. in Hinder et al. (2018), is commonly used in such procedure since this scalar quantity defines the optimal matched filter in the GW detection problem. The mismatch can either play the role of the LF to minimize (e.g. Joshi et al., 2023) or can be used as a validation metric after minimization of another LF. For instance, the authors of Habib and Huerta (2019) computed their LFs from the evolution of the dimensionless quantity $M\omega$, with $M$ the BBH total mass and $\omega$ the mean orbital frequency, and then calculated the mismatch to quantify the accuracy of the characterization of the eccentric NR simulation at hand.
@@
Alternatively, one can choose the $L_{2}$-norm, the distance between vectors of strain data in time domain as the LF to compare a model (waveform, surrogate model to train) from a set of NR simulations Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024b); Rink et al. (2024). We will thus also include this $L_{2}$-norm in our analysis.
@@ III.1 NR simulations and model waveform
For this study, we use $10$ eccentric simulations, as listed in Table 2, from the Simulating eXtreme Spacetimes (SXS) catalog (Boyle et al., 2019; SXS Collaboration, ; SXS, ) and run by Islam et al. (2021). We sort them in this table by $e_{\text{gw}}$ at $18$Hz from the measurements of Fig. 1. We use the implementation in the Python library gw_eccentricity  [^2] https://github.com/vijayvarma392/gw_eccentricity from Shaikh et al. (2023) to compute $e_{\text{gw}}$ in this work. We explain the chosen method later in Section III.2.
@@
To assign $e^{\text{best-fit}}$ to these simulations, we use the publicly available eccentric waveform model TEOBResumS-Dalí  [^3] https://bitbucket.org/teobresums/teobresums/. We have used the git commit 0f19532a9a1012b556bd4c72da10a1317d0e8751. To compute $e_{\text{gw}}$ and the LF $\mathcal{E}$, we will make use of the dominant (2, 2) mode only and output physical units strain from TEOBResumS-Dalí, In addition, since TEOBResumS-Dalí allows to define the reference frequency as the waveform frequency at apastron, periastron or the mean of the two, we choose the latter definition as in Knee et al. (2022). The inclusion of higher modes has notably been investigated by e.g. Habib and Huerta (2019); Bonino et al. (2023), and is beyond the scope of our study. It should also be kept in mind that the definition of the eccentric initial conditions are sensitive to changes in input parameters $e_{0},f_{0}\sim\mathcal{O}(10^{-6})$ Nagar and Rettegno (2021).
@@ III.2 Measuring $e_{\text{gw}}$
Figure 1: Left: Frequency evolution of eccentricity $e_{\text{gw}}$ in the simulations from Table. 2. Right: $e_{\text{gw}}$ ratio between the ResidualAmplitude and Amplitude method from gw_eccentricity Shaikh et al. (2023). The former is more robust to the small eccentric NR strains in our set, justifying our preference. We resort to physical units by choosing a total binary mass $M=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc.
@@ III.3.2 Loss Functions
    We will note, for simplicity, $\mathcal{E}_{\mathpzc{h}_{22}}(f_{0},e_{0})=\|\delta\mathpzc{h}_{22}\|$, also emphasizing that this scalar is a distance in time-domain functions and not a correlation measure of Fourier space vectors as the mismatch. This LF is at the core of works such as Blackman et al. (2017a); Islam et al. (2021, 2022); Islam (2024b); Rink et al. (2024), notably to fit surrogate parametric models to reference NR simulations.
@@ V Conclusion and remarks
In addition, the recent study Bonino et al. (2024) would corroborate this prescription on the mismatch, although not in the context of numerical searches as in our work. The authors, after assigning eccentricities to their own NR simulations set through a mapping procedure of the initial conditions, find $\mathcal{M}\sim 10^{-3}-10^{-2}$ (Figure 12 of current version) as a robustness metric rather than a LF. We argue that $C_{\lambda}$ is merely an alternate LF to deal with both parametric model limitations and mismatch requirements. A simple to implement LF like $C_{\lambda}$ will be particularly useful in the context of fitting surrogates to reference strains, among the practical applications that motivated our study. We leave new developments on fitting procedures invoking $e_{\text{gw}}$ in addition to classic LFs, with numerical searches or other strategies as in Bonino et al. (2024), to future studies.
@@ A.2 LF minimization results
Table 2: Characterization of SXS simulations with TEOBResumS-Dalí using different LFs, along with standardized $e_{\text{gw}}$ measurements. Blue and red entries correspond to Figure 11. The leftmost column indicates the chosen LF, and the "Values" are computed from the TEOBResumS-Dalí waveform that minimizes this LF. From left to right, the SXS:BBH:XXXX IDs are sorted by increasing $e_{\text{gw}}$ from Table 2. Reference frequencies $f_{0}$ in Hz and $f_{m}=18$Hz correspond to total mass $M=m_{1}+m_{2}=60M_{\odot}$ and luminosity distance $D_{L}=400$Mpc. Cases where the best-fit model has $\delta e_{\text{gw}}\sim 100\%$ correspond to a preferred quasi-circular model by that LF, which occurs when $LF\sim\mathcal{M}$ for the small eccentricity NR simulations.